OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. P(6, 2), Q(8, -4) Explain it too please :)

OpenStudy (anonymous):

I found out how to do it if anyone has the question. So Standard form is Ax+By=C. but we cant put it like that right away. First, we have to find the slope. The slope would be 1 in this case. after you find the slope you put it in point slope form, which is (y-y1)=m(x-x1). so our equation needs one point from he question which i am using p. so (y-2)=1(x-6) After you simplify this, you get y-2=x-6. Add 2 to the right side, and subtract x from the left side. after you do that, the equation should be -x+y=-4. but A can not be negative in this equation. so if a is -1, multiply every term by -1, which changes each of the operations in the equation to change. So the answer to the question would be x-y=4! Hope this helps some of you!.

OpenStudy (anonymous):

leaving the queston open in case some people find a mistake, but it should be right.